OpenStudy (anonymous):

Assume the domain of f and f^-1 is (-infinity,infinity). Solve the equation for t: f((1-2t)/(1+2t)) = 7 ; f^-1 (7) =-3

OpenStudy (anonymous):

\[ f^{-1}(f(x)) =x \]

OpenStudy (anonymous):

Let \[ x=\frac{1-2t}{1+2t} \]

OpenStudy (anonymous):

\[ f(x) = 7 \]take \(f^{-1}\) of both sides and get: \[ x = f^{-1}(7) = -3 \]

OpenStudy (anonymous):

And so \[ x=-3\implies \frac{1-2t}{1+2t} = -3 \]

OpenStudy (anonymous):

Can you solve for \(t\) now?

OpenStudy (anonymous):

Yes thank you!